# Unsupervised Attributed Dynamic Network Embedding with Stability Guarantees

### v1

@@ D.2 Method comparison
The following list shows unsupervised dynamic attributes methods we didn’t include in our method comparison and the reason why.
@@
  • [Xu et al., 2020a] - First, the code doesn’t allow for time-varying covariates. Second, it is extremely time-consuming for data sets with large $p$; for example, one epoch on one batch for DBLP takes more than 10 min, meaning the whole computation would be more than 1000 hours. The authors presented examples with a maximum $p=200$, while for the data sets considered in the paper, $p$ is at least 4000.
@@ D.4 Additional tables
Table 3: AUC of link prediction on AUASE embeddings for varying $\alpha$.

### v2

@@ D.1 Experimental setup details
ogbn-mag. The raw data can be downloaded from https://ogb.stanford.edu/docs/nodeprop/. The data spans years from 2010 to 2019, each year is a time point. We consider authors which published at least ten paper in the whole time frame.
To construct the adjacency matrices, we form an edge if two authors collaborated on at least one paper within that time point. Each paper is associated with a 128-dimensional word2vec feature vector, to construct the attributes matrices for each author we average the feature vectors of the papers they published in that time point. An author label is set to the label associated with the majority of the papers they published in that time point.
We use ogbn-mag data to construct a datasets which is suited to our dynamic tasks. The data format and the tasks are at author-level, different than the usual task performed on obgn-mag (‘the task is to predict the venue of each paper, given its content, references, authors, and authors’ affiliations’). That is because we are interested in dynamic tasks; authors are dynamic nodes, unlike papers, and their attributes and labels change over time. This creates a much harder task, hence our performance cannot be directly compared to other performances on obgn-mag to predict papers labels.
CONN, Glodyne, DySAT and DyRep are not computationally efficient enough to compute the embeddings of ogbn-mag. We report here the computational barriers for each methods.
  • CONN: the authors’ code is designed with dense matrices, which would require about 2T of memory for ogbn-mag matrices.
  • Glodyne: we capped computational time after 48h.
  • Dyrep: it requires an unreasonably large amount of memory. The largest dataset the authors’ analyses has only $\approx 10000$ nodes.
@@ D.2 Method comparison
The following list shows unsupervised dynamic attributes methods we did not include in our method comparison and the reason why.
@@
  • [Xu et al., 2020a] - First, the code does not allow for time-varying covariates. Second, it is extremely time-consuming for data sets with large $p$; for example, one epoch on one batch for DBLP takes more than 10 min, meaning the whole computation would be more than 1000 hours. The authors presented examples with a maximum $p=200$, while for the data sets considered in the paper, $p$ is at least 4000.
@@ D.4 Additional tables
Table 3: AUC of link prediction on AUASE embeddings for varying $\alpha$. For $\alpha=1$ the method is using only the attributes and not the network, contrarily to AUASE for $\alpha\in(0,1).$